### v1

# Severity Prediction in Mental Health: LLM-based Creation, Analysis, Evaluation of a Novel Multilingual Dataset

Abstract
Large Language Models (LLMs) are increasingly being integrated into various medical fields, including mental health support systems. However, there is a gap in research regarding the effectiveness of LLMs in non-English mental health support applications. To address this problem, we present a novel multilingual adaptation of widely-used mental health datasets, translated from English into six languages (Greek, Turkish, French, Portuguese, German, and Finnish). This dataset enables a comprehensive evaluation of LLM performance in detecting mental health conditions and assessing their severity across multiple languages. By experimenting with GPT and Llama, we observe considerable variability in performance across languages, despite being evaluated on the same translated dataset. This inconsistency underscores the complexities inherent in multilingual mental health support, where language-specific nuances and mental health data coverage can effect the accuracy of the models. Through comprehensive error analysis, we emphasize the risks of relying exclusively on large language models (LLMs) in medical settings (e.g., their potential to contribute to misdiagnoses). Moreover, our proposed approach offers significant cost savings for multilingual tasks, presenting a major advantage for broad-scale implementation.
Severity Prediction in Mental Health: LLM-based Creation, Analysis, Evaluation of a Novel Multilingual Dataset
Konstantinos Skianis1, John Pavlopoulos2, A. Seza Doğruöz3 1University of Ioannina, Greece 2Athens University of Economics and Business, Greece 3Universiteit Gent, Belgium Corresponding author: kskianis@cse.uoi.gr
1 Introduction
Large language models (LLMs) (e.g., GPT-3, BERT, and their multilingual variants) are trained on vast amounts of text data and demonstrate an impressive ability to generate human-like text, perform complex language tasks Devlin et al. (2019); Brown et al. (2020). One of the most promising applications of LLMs is in the healthcare domain, where they have the potential to assist medical specialists with the detection, diagnosis, and treatment of health conditions.
Mental health disorders such as depression, anxiety, and post-traumatic stress disorder (PTSD) are widespread across the globe, presenting major public health challenges World Health Organization (2021). Traditional diagnostic approaches typically involve self-reported questionnaires, clinical interviews, and standardized assessments performed by trained professionals Kessler and Üstün (2004). However, these processes are also time-consuming and labor intensive which leads to delays in treatment and increased waiting times for the patients.
Therefore, leveraging large language models (LLMs) to detect mental health symptoms from textual data presents a promising alternative. These models have the ability to process vast amounts of textual content (e.g., social media posts, online forum discussions, personal narratives), and detect specific linguistic indicators related to mental health issuesGuntuku et al. (2019); Chancellor et al. (2019). In this way, health professionals can be assisted with the early detection and intervention of mental health symptoms, reduce delays for patients whose symptoms might otherwise go unnoticed.
However, most studies about applying LLMs into mental health research are still in English and there is a need to explore this research area for other languages Raihan et al. (2024).
Our study fills these gaps with the following research questions:
  • How accurately can LLMs predict the severity of mental health conditions in English-language through user-generated content?
  • When English-language mental health datasets are translated into other languages using LLMs, does the model maintain similar accuracy in predicting the severity of mental health conditions across these languages?
We answer these research questions with the following key contributions:
  • We create a novel multilingual dataset (covering English, Turkish, French, Portuguese, German, Greek, and Finnish) derived from user-generated social media content, which is automatically translated from English using an LLM. [^1] The dataset is publicly available at the following link: https://github.com/y3nk0/multilingual-mental-severity-prediction.
  • We perform an extensive evaluation of the effectiveness of an LLM in predicting the severity of mental health conditions across these different languages.
  • We provide an in-depth analysis of the model’s performance in each language, offering insights into how linguistic diversity affects the task and dataset-specific outcomes.
To the best of our knowledge, this is the first study to utilize LLMs for creating and evaluating multilingual adaptations of existing mental health NLP datasets, specifically designed to detect the severity of mental health conditions. Our study pushes forward the creation of AI-powered tools and datasets for mental health diagnosis that are both inclusive and adaptable, emphasizing the critical need for approaches that account for cultural and linguistic diversity in mental health care, especially in non-English contexts with cost-efficient methods.
2 Related work
Early investigations into the use of NLP for mental health concentrated on identifying specific symptoms from textual data through machine learning methods. This often involved analyzing social media content, online health forum discussions, and other text-based sources to find indicators of mental health conditions like depression, anxiety and others De Choudhury and De (2014); Coppersmith et al. (2014); De Choudhury and De (2014); Coppersmith et al. (2015); Shing et al. (2018); Guntuku et al. (2019). Traditional approaches primarily relied on machine learning algorithms with hand-crafted features, which lacked the sophistication of current LLMs.
Multilingual Capabilities of LLMs
Although training and evaluation details are not always fully disclosed, LLMs are assumed to be capable of processing and generating text in multiple languages. For instance, the XLM-R model (trained on extensive multilingual data), has demonstrated strong cross-lingual performance. As noted by Conneau et al. (2020), XLM-R surpasses earlier models on various tasks, highlighting the benefits of leveraging large-scale multilingual data.
There are also examples of LLMs excelling in practical applications (e.g., translation and multilingual question answering). For instance, a model created by Xue et al. (2021), is pre-trained on a diverse set of languages and demonstrates strong capabilities in translating and answering questions in multiple languages. The model’s architecture and training methods allow it to generalize across languages effectively, providing robust performance even in languages with limited training data.
Figure 1: An illustration of our proposed methodology.
However, major obstacles still hinder the achievement of truly equitable performance across all languages Zhang et al. (2020). One of the primary challenges lies in low-resource languages, which suffer from a scarcity of digitally available text data. This often leads to under-performance by LLMs, producing less reliable outputs for such languages. Tackling this issue requires not only more inclusive and comprehensive data collection but also innovative approaches in transfer learning that can better leverage sparse resources Doğruöz and Sitaram (2022).
Moreover, language is intricately tied to cultural and societal frameworks, making it particularly difficult to capture the subtle, context-driven nuances that vary from one culture to another. This adds another layer of complexity to multilingual NLP tasks. Addressing these intricacies demands a twofold approach: developing more advanced, culturally-aware algorithms and incorporating datasets that are truly representative of linguistic and cultural diversity Doğruöz et al. (2023). Without these efforts, the full potential of multilingual LLMs will remain inaccessible to many, limiting their impact on global applications.
LLMs for Mental Health
LLMs excel at detecting mental health symptoms with greater precision by leveraging their advanced understanding of context and semantics. Notable examples include BioBERT Lee et al. (2020) and ClinicalBERT Huang et al. (2019), which have been pre-trained on biomedical literature and clinical notes. Additionally, models like MentalBERT Ji et al. (2022), DisorBERT Aragon et al. (2023), SuicidalBERT Anonymous , MentalLLM Xu et al. (2024) and MentalLlama Yang et al. (2023) are specifically trained on mental health-related social media data. Nevertheless, using LLMs for detecting mental health issues may introduce a number of advantages, as well as disadvantages De Choudhury et al. (2023). Additionally exploiting social network posts represents a promising frontier in both artificial intelligence and mental health care Harrigian et al. (2020). LLMs (e.g., GPT-4 and BERT) can analyze vast amounts of textual data from social media platforms automatically to identify linguistic patterns associated with mental health conditions such as depression, anxiety, and suicidal ideation Ji et al. (2023).
Recently, a wide range of social network datasets focusing on mental health have become accessible Raihan et al. (2024). The authors compiled posts from platforms like Reddit and Twitter that pertain to conditions such as depression, PTSD, schizophrenia, and eating disorders. In addition, they fine-tuned several models using smaller, publicly available annotated mental health datasets to facilitate labeling for their newly introduced MentalHelp dataset. However, the dataset is limited to English-language posts, which constrains its application and hinders further research in multilingual contexts.
3 Proposed Methodology
We present a comprehensive methodology for leveraging LLMs and social media posts (in English), in order to translate them into multiple languages and predict mental health conditions accordingly.
First, we translate the social media posts across the proposed languages via an LLM. Then, these translations are fed into a prompt that asks the LLM to predict specific severity levels of mental health conditions. Finally, the predicted classes are compared with the ground truth labels and the evaluation metrics are computed.
Our goal is to propose a methodology that can be used to evaluate LLMs on multilingual detection of mental health conditions in languages where minimal relative linguistic resources are publicly available. An illustration of the proposed approach for evaluating LLMs for multilingual detection of mental health conditions is shown in Figure 1.
Dataset Reference Category #Classes #Instances Labels (#Support)
Dep-Severity Naseem et al. (2022) Depression 4 3553 Minimum (2587)
Mild (290)
Moderate (394)
Severe (282)
Sui-Twi Mirza Ibtihaj et al. (2020) Suicide 3 1300 Not Suicidal (515)
Suicidal Ideation (455)
Potential Suicide Attempt (330)
Table 1: Benchmark datasets used in our study.
4 Experiments
4.1 Datasets
In order to test the proposed methodology we are using the following datasets:
Dep-Severity
Introduced by Naseem et al. (2022), this dataset comprises posts from the social media platform Reddit, specifically addressing the varying levels of depression. These posts, all in English, are pre-labeled according to four severity levels: minimal, mild, moderate, and severe depression. Notably, the dataset is highly imbalanced, with the majority of posts classified under the minimal severity level.
Sui-Twi
This dataset (Mirza Ibtihaj et al., 2020) contains texts from Twitter (in English), including 1,300 suicidal ideation tweets. The authors gathered posts that contained the word ‘suicide’ and labeled them manually into three classes ”not suicidal", ”indicative of suicidal ideation", and ”indicative of a potential suicide attempt". In this dataset, the classes are balanced in terms of frequency of occurence.
We deliberately chose multi-class datasets to increase the difficulty for the LLM, as binary classification tasks would have been simpler to solve. Table 1 provides a summary of key statistics for the two datasets utilized in our study.
Dataset Prediction prompt
Dep-Severity “Categorize the following text with 1 of the 4 depression severity levels (0: Minimum, 1: Mild, 2: Moderate, 3: Severe)"
Sui-Twi “Categorize to 0 if the text is not suicidal, to 1 if the text is almost related to suicidal ideation, and to 2 if the text indicates a potentially suicidal user"
Table 2: The used prediction prompts.
4.2 Settings
We use GPT3.5-turbo Brown et al. (2020), GPT-4o-mini via the OpenAI API and Llama3.1 locally to translate the posts and predict the labels. Recently introduced models, such as MentaLLaMA Yang et al. (2024), primarily focus on English, making them unsuitable for use in multilingual settings.The temperature parameter is set to 0, so the outcome is reproducible (regarding the translations and predictions). We experiment with no examples in the prompt (0-shot) and one example per class (1-shot). The task is solved as a classification problem, where we compare the predicted classes with the ground-truth ones, reporting Precision (Pr), Recall (Rec) and F1. We experiment with seven languages from major language families: Turkish (Altaic), French, Portuguese (Romance), English, German (Germanic), Finnish (Uralic), Greek (Hellenic). Table 2 shows the used prompts for predicting the severity levels.
English Turkish French Portuguese German Greek Finnish
Class Pr Rec F1 Pr Rec F1 Pr Rec F1 Pr Rec F1 Pr Rec F1 Pr Rec F1 Pr Rec F1
minimum 0.98 0.14 0.25 0.93 0.03 0.05 0.99 0.15 0.27 0.98 0.11 0.20 0.97 0.24 0.38 0.99 0.07 0.14 0.98 0.08 0.15
mild 0.04 0.15 0.07 0.02 0.01 0.01 0.05 0.17 0.08 0.04 0.07 0.05 0.08 0.32 0.12 0.04 0.17 0.06 0.05 0.19 0.08
moderate 0.13 0.22 0.17 0.10 0.62 0.17 0.15 0.39 0.21 0.12 0.28 0.17 0.16 0.42 0.23 0.14 0.55 0.23 0.12 0.31 0.17
severe 0.13 0.71 0.22 0.13 0.39 0.19 0.15 0.58 0.23 0.12 0.74 0.20 0.16 0.35 0.22 0.16 0.28 0.20 0.12 0.55 0.20
Macro avg 0.32 0.30 0.17 0.29 0.26 0.11 0.33 0.32 0.20 0.31 0.30 0.16 0.34 0.33 0.24 0.33 0.27 0.16 0.32 0.28 0.15
Table 3: GPT-3.5 with 0-shot learning on Dep-Severity, measuring Precision, Recall, and F1 per class per language. The last row shows the macro averages. The best F1 per class is shown in bold. The best macro-averaged F1 across languages is underlined.
English Turkish French Portuguese German Greek Finnish
Class Pr Rec F1 Pr Rec F1 Pr Rec F1 Pr Rec F1 Pr Rec F1 Pr Rec F1 Pr Rec F1
not suicidal 0.78 0.53 0.63 0.58 0.44 0.50 0.58 0.66 0.62 0.63 0.53 0.58 0.56 0.67 0.61 0.58 0.67 0.63 0.57 0.52 0.55
ideation 0.46 0.78 0.58 0.36 0.70 0.47 0.43 0.40 0.42 0.43 0.57 0.49 0.39 0.42 0.41 0.49 0.50 0.49 0.35 0.42 0.38
potential 0.69 0.36 0.47 0.54 0.04 0.07 0.54 0.47 0.50 0.54 0.41 0.47 0.51 0.31 0.39 0.65 0.47 0.55 0.49 0.39 0.43
Macro avg 0.64 0.56 0.56 0.49 0.39 0.35 0.51 0.51 0.51 0.53 0.51 0.51 0.49 0.47 0.47 0.57 0.55 0.55 0.47 0.45 0.45
Table 4: GPT-3.5 with 0-shot learning on Sui-Twi, measuring Precision, Recall, and F1 per class per language. The last row shows the macro averages. The best F1 per class is shown in bold. The best macro-averaged F1 across languages is underlined.
English Turkish French Portuguese German Greek Finnish
Class Pr Rec F1 Pr Rec F1 Pr Rec F1 Pr Rec F1 Pr Rec F1 Pr Rec F1 Pr Rec F1
minimum 0.97 0.32 0.48 0.97 0.18 0.30 0.97 0.30 0.46 0.97 0.31 0.47 0.96 0.36 0.52 0.98 0.26 0.41 0.97 0.26 0.40
mild 0.07 0.26 0.11 0.04 0.18 0.07 0.05 0.14 0.07 0.07 0.22 0.11 0.06 0.14 0.08 0.05 0.13 0.07 0.05 0.17 0.07
moderate 0.18 0.41 0.25 0.15 0.42 0.23 0.16 0.50 0.24 0.16 0.41 0.23 0.15 0.42 0.22 0.15 0.52 0.23 0.15 0.48 0.23
severe 0.19 0.43 0.26 0.17 0.48 0.25 0.18 0.44 0.26 0.17 0.49 0.25 0.18 0.48 0.26 0.17 0.40 0.24 0.18 0.40 0.25
Macro avg 0.35 0.35 0.27 0.33 0.31 0.21 0.34 0.34 0.26 0.34 0.36 0.27 0.34 0.35 0.27 0.34 0.33 0.24 0.34 0.33 0.24
Table 5: GPT-4o-mini with 0-shot learning on Dep-Severity, measuring Precision, Recall, and F1 per class per language. The last row shows the macro averages. The best F1 per class is shown in bold. The best macro-averaged F1 across languages is underlined.
English Turkish French Portuguese German Greek Finnish
Class Pr Rec F1 Pr Rec F1 Pr Rec F1 Pr Rec F1 Pr Rec F1 Pr Rec F1 Pr Rec F1
not suicidal 0.62 0.88 0.73 0.83 0.61 0.70 0.67 0.80 0.73 0.66 0.71 0.68 0.62 0.87 0.72 0.64 0.81 0.71 0.62 0.84 0.71
ideation 0.48 0.32 0.38 0.46 0.60 0.52 0.51 0.49 0.50 0.43 0.55 0.49 0.42 0.38 0.40 0.56 0.44 0.49 0.41 0.32 0.36
potential 0.75 0.61 0.67 0.45 0.45 0.45 0.78 0.60 0.68 0.79 0.41 0.54 0.78 0.38 0.52 0.81 0.73 0.77 0.69 0.51 0.58
Macro avg 0.62 0.60 0.59 0.58 0.55 0.56 0.66 0.63 0.64 0.63 0.55 0.57 0.61 0.54 0.55 0.67 0.66 0.66 0.57 0.56 0.55
Table 6: GPT-4o-mini with 0-shot learning on Sui-Twi, measuring Precision, Recall, and F1 per class per language. The last row shows the macro averages. The best F1 per class is shown in bold. The best average F1 across languages is underlined.
English French Portuguese German
Class Pr Rec F1 Pr Rec F1 Pr Rec F1 Pr Rec F1
minimum 0.88 0.12 0.22 0.91 0.23 0.37 0.95 0.18 0.30 0.95 0.25 0.40
mild 0.04 0.20 0.07 0.04 0.13 0.07 0.05 0.14 0.08 0.05 0.15 0.08
moderate 0.15 0.60 0.25 0.16 0.58 0.24 0.14 0.76 0.24 0.14 0.41 0.21
severe 0.18 0.22 0.20 0.12 0.26 0.17 0.20 0.11 0.14 0.13 0.41 0.20
Macro avg 0.32 0.29 0.19 0.31 0.30 0.21 0.34 0.30 0.19 0.32 0.31 0.22
Table 7: Llama3.1 with 0-shot learning on Dep-Severity, measuring Precision, Recall, and F1 per class per language. The last row shows the macro averages. The best F1 per class is shown in bold. The best average F1 across languages is underlined.
English French Portuguese German
Class Pr Rec F1 Pr Rec F1 Pr Rec F1 Pr Rec F1
not suicidal 0.76 0.54 0.63 0.44 0.50 0.47 0.67 0.58 0.62 0.49 0.63 0.55
ideation 0.43 0.40 0.41 0.38 0.15 0.21 0.47 0.50 0.49 0.47 0.25 0.33
potential 0.45 0.68 0.54 0.23 0.37 0.28 0.48 0.53 0.50 0.39 0.45 0.42
Macro avg 0.55 0.54 0.53 0.35 0.34 0.32 0.54 0.54 0.54 0.45 0.45 0.43
Table 8: Llama3.1 with 0-shot learning on Sui-Twi, measuring Precision, Recall, and F1 per class per language. The last row shows the macro averages. The best F1 per class is shown in bold. The best average F1 across languages is underlined.
English Turkish French Portuguese German Greek Finnish
Class Pr Rec F1 Pr Rec F1 Pr Rec F1 Pr Rec F1 Pr Rec F1 Pr Rec F1 Pr Rec F1
minimum 0.90 0.14 0.25 0.76 0.23 0.36 0.91 0.11 0.19 0.90 0.11 0.20 0.76 0.75 0.76 0.79 0.02 0.05 0.68 0.16 0.25
mild 0.03 0.06 0.04 0.07 0.10 0.08 0.04 0.04 0.04 0.05 0.14 0.07 0.07 0.08 0.07 0.05 0.10 0.06 0.03 0.08 0.05
moderate 0.14 0.65 0.23 0.11 0.29 0.16 0.13 0.44 0.20 0.15 0.51 0.23 0.16 0.23 0.19 0.14 0.26 0.18 0.13 0.39 0.19
severe 0.15 0.35 0.21 0.09 0.40 0.14 0.11 0.61 0.19 0.13 0.48 0.20 0.16 0.07 0.09 0.10 0.73 0.17 0.10 0.37 0.16
Macro avg 0.30 0.30 0.18 0.26 0.26 0.18 0.30 0.30 0.16 0.31 0.31 0.18 0.29 0.28 0.28 0.27 0.28 0.12 0.23 0.25 0.16
Table 9: GPT-3.5 with 1-shot learning on Dep-Severity by measuring Precision, Recall, and F1 per class per language. The last row shows macro averages for precision, recall, and F1. Best F1 per class is shown in bold. Best average F1 across languages is underlined.
English Turkish French Portuguese German Greek Finnish
Class Pr Rec F1 Pr Rec F1 Pr Rec F1 Pr Rec F1 Pr Rec F1 Pr Rec F1 Pr Rec F1
not suicidal 0.72 0.11 0.20 0.55 0.26 0.35 0.47 0.03 0.05 0.59 0.21 0.31 0.50 0.16 0.25 0.72 0.03 0.07 0.39 0.09 0.14
ideation 0.37 0.81 0.51 0.36 0.25 0.29 0.68 0.09 0.16 0.37 0.36 0.37 0.51 0.33 0.40 0.17 0.00 0.00 0.34 0.10 0.15
potential 0.39 0.26 0.31 0.36 0.80 0.50 0.27 0.98 0.42 0.30 0.61 0.40 0.32 0.82 0.46 0.26 0.99 0.41 0.27 0.87 0.42
Macro avg 0.49 0.40 0.34 0.42 0.44 0.38 0.47 0.37 0.21 0.42 0.39 0.36 0.44 0.44 0.37 0.38 0.34 0.16 0.33 0.35 0.24
Table 10: GPT-3.5 with 1-shot learning on Sui-Twi by measuring Precision, Recall, and F1 per class per language. The last row shows macro averaged scores. The best F1 per class is shown in bold and the best average F1 across languages is underlined.
4.3 Preliminary Prompting
Before presenting any posts to our LLM, whether for translation or classification, we first inquired how it would approach categorizing them into varying levels of depression severity. The LLM explained that it would begin by identifying specific language cues linked to depressive states, such as:
  • Recurrent negative emotions like sadness or a sense of hopelessness.
  • Self-critical thoughts or feelings of worthlessness.
  • Expressions reflecting isolation or withdrawal from social interactions.
  • Observable shifts in behavior or routines, such as disrupted sleep or changes in appetite.
  • Mentions of emotional suffering or psychological distress.
Furthermore, the LLM would attempt to adapt these signals to the four depression severity levels in the context of social media content. For instance:
  • Level 1 (Minimal): Posts that show infrequent or minor expressions of sadness or frustration.
  • Level 2 (Mild): Posts where negative emotions or notable behavioral shifts are more consistent.
  • Level 3 (Moderate): Posts indicating substantial interference with daily life or clear markers of emotional distress.
  • Level 4 (Severe): Posts showing extreme emotional distress, risk of self-harm, or complete social disengagement.
From this, we can infer that the LLM anticipates posts containing broad indicators of depressive symptoms.
Similarly, for the second dataset, we expect the LLM to detect various linguistic and contextual features in the social media posts to classify them. Some key aspects that it considers per class are:
  1. 1.
    Not Suicidal:
    • General Tone: Neutral, positive, or typical negative expressions not related to self-harm or life-threatening situations.
    • Content: Discussions about daily activities, hobbies, interests, work, or social interactions without any reference to self-harm or death.
    • Language: Absence of words or phrases related to self-harm, death, hopelessness, or helplessness.
  2. 2.
    Indicative of Suicidal Ideation:
    • Expression of Hopelessness: Phrases indicating a lack of hope for the future, such as “I can’t see any way out", “Things will never get better", or "There’s no point in trying".
    • Self-Hatred or Worthlessness: Statements like “I’m a burden", “Everyone would be better off without me," or "I hate myself".
    • Mentions of Death or Dying: Expressions about wanting to die or not wanting to live, such as “I wish I were dead", “I want to end it all", or “Life isn’t worth living".
    • Social Withdrawal: Indications of isolating oneself from others, such as “I don’t want to see anyone", or “I prefer to be alone all the time".
  3. 3.
    Indicative of a Potential Suicide Attempt:
    • Detailed Plans or Intentions: Descriptions of specific methods or plans to commit suicide, such as “I have pills ready," “I’m going to jump off a bridge," or “I’ve written my goodbye letter."
    • Goodbye Messages: Posts that read like final farewells, such as “Goodbye everyone, it’s been a pleasure," or “I’m sorry for everything, I won’t be around much longer."
    • Previous Attempts: References to past suicide attempts, particularly if combined with current suicidal thoughts, such as “I’ve tried before and failed, but not this time."
    • Lethal Means: Mentions of acquiring or possessing means to commit suicide, like “I bought a gun," “I have enough pills," or “I found a rope."
When asked, the LLM presents a detailed description of the three classes, which seems promising.
4.4 Experimental results
The results are organized and presented in the following subsections: classification in the source language, classification in the target language, classification across languages, and further analysis of LLMs performance with multilingual examples.
Source language
Initially, we experimented with the data in their source language (English), to set the baseline performance. That is, no translation step has been performed at this stage. When using GPT-3.5, the first column in Tables 3 and 4 presents the results on depression and suicide detection tasks respectively. The best F1 in both tasks is achieved for the lowest severity/indication. In Dep-Severity, the best performance is achieved for the two edges of lowest (F1=0.48) and highest severity (F1=0.26). In Sui-Twi, on the other hand, F1 drops as the suicidal likelihood increases, from 0.73 to 0.67. The task of Dep-Severity is more challenging (F1=0.17) compared to Sui-Twi (F1=0.56). When using the superior GPT-4o-mini (Tables 6 and 5), F1 increases to 0.27 and to 0.59, respectively. This can be attributed to the fact that detecting specific levels of depression is harder than in the case of suicidal ideation, as depression severity levels are considered less distinct. Therefore, it is probably more challenging for the LLM to distinguish.
Target languages
GPT-3.5 has a varying performance when operating on content translated from English to target languages. In Dep-Severity (Table 3), performance drops in Turkish (-6), Portuguese (-1), Greek (-1), and Finish (-2). On the other hand, the performance increases in French (+3) and German (+7). In Sui-Twi (Table 3), performance drops in all target languages, with the drop for Turkish (-17) being the most considerable one. This can be explained by the limited number of Turkish resources in health, especially in the domain of mental health Çöltekin et al. (2023).
GPT-4o-mini in Dep-Severity (Table 5) performs equally well in Portuguese and German (as in English), but this is not true for the rest of the target languages. In Sui-Twi, performance drops for most target languages. However, we observe again the phenomenon of better performance for a target language, namely for French (+5) and Greek (+7), as was the case for GPT-3.5 in French and German. Therefore, we find that by translating to French, performance increases for both models. The best performance is found for Greek (F1=0.66) and the worst for German and Finnish (-4).
Classification across languages
Figures 2 and 3 present F1 in boxplots, summarising the performance of GPT-3.5 across languages per task. In Dep-Severity, the lowest (minimum) severity class is characterised by a wide range, from 0.15 to 0.25. The whiskers are also wide, from 0.05 (Turkish) to 0.38 (German). Mild severity is mishandled across languages while for the classes of moderate and severe, the model achieves a relatively high score across languages.
Figure 2: F1 of GPT-3.5 across languages (vertically) per class (box) for Dep-Severity.
Figure 3: F1 of GPT-3.5 across languages (vertically) per class (box) for Sui-Twi.
In Sui-Twi, the best performance is achieved for the “non-suicidal" category. This is also the major class while slightly worse performance is achieved for the other two classes. For all three classes the performance is similarly distributed across languages.
From zero to single shot
Tables 9 and  10 present the results when 1-shot learning is employed. In English, performance drops considerably in Sui-Twi and remains similarly low in Dep-Severity. Again, performance is better for a target language, compared to the performance for the source language, but considerably lower (or equal) than what achieved with zero-shot learning. Better performance could be observed if more examples were provided, but this would also increase the cost of the experiments by approx. doubling the size of the prompt.
Experimenting with Llama-3.1
Tables 7 and 8 show the performance of Llama-3.1 in both tasks, using 0-shot, for the languages it was available for. Although it is better in English compared to GPT-3.5 (both 0- and 1-shot), it is slightly worse in target languages. We observe high variability in Sui-Twi, with French being the worst and Portuguese the best, but not in Dep-Severity. In both tasks, the best performance is not achieved in English.
Further analysis with examples
Figure 4: An example of translation (English to Greek) from the suicide dataset. The original words and their translations are colored similarly.
Lastly, Figure 4 presents two instances of the suicide dataset. Regarding the first example, while the translation from English to Greek appears to be correct in terms of conveying the same message, the English text is classified as "Potential suicide attempt", while the Greek text is classified as "Suicidal ideation". The LLM might be interpreting the context differently due to nuances in the languages. In English, "attempted suicide" explicitly suggests a recent suicide attempt, which clearly falls under "Potential suicide attempt." In Greek, the model might be weighing other parts of the text more heavily, leading to a classification of "Suicidal ideation." This could be due to differences in the training data, language model intricacies, or contextual emphasis.
For the second example, we observe that the translation is accurate and conveys the same message as the original English text. Both the English and Greek texts are classified consistently and correctly as "suicidal ideation".
4.5 Resources and cost of experiments
The total cost for running the GPT3.5 experiments via its API was less than \$30, showing the minimal resources required to conduct them, without the need for expensive GPU infrastructure or fine-tuning. For example, we experimented with MentaLLaMA-chat-7B Yang et al. (2024), which is a 27GB model. Its size makes it prohibitive when limited resources are available. Eventually we ended up using the quantized version, which did not output meaningful results. Our methodology for utilizing resources is potentially promising for extending medical data sets in English into other languages as well.
5 Discussion
Despite the hype of using LLMs in mental healthcare for commercial applications, there is an increasing need for more research to understand the pitfalls and dangers of using LLMs across languages, diseases, and patient populations for mental health care. Considering the variation in the performance of the LLM across languages, there is a need for utmost precautions not to rely on LLMs solely in any healthcare setting. As stated by Stade et al. (2024), mental health diagnosis should never be left to automatic systems only and should never replace human professionals due to possible errors and/or misdiagnosis illustrated above.
LLMs exhibit considerable variability in performance across languages, despite being evaluated on the same translated dataset. This inconsistency underscores the complexities inherent in multilingual mental health support, where language-specific nuances and mental health data coverage can effect the accuracy of the models. We observe performance gains in both tasks with stronger models and by translating to a specific target language (e.g., in German, for depression severity) before detection. On the other hand, we also observe that translating to Turkish yields the worst performance in both tasks, which could be explained by the limited number of Turkish resources in health, especially in the domain of mental health Çöltekin et al. (2023).
6 Conclusion
We presented a novel multilingual dataset designed to evaluate the ability of LLMs to predict the severity of mental health conditions in user-generated posts. Using English as the source language, we translated the posts to French, German, Turkish, Greek, Finnish, and Portuguese, using LLM-based machine translation. Then, we prompted the LLM to predict the severity of two mental health conditions across all languages. Our findings show that the LLM performance varies depending on the language. Also, the best performance is not necessarily achieved in the source (English) language. Although not tested explicitly in low resource languages, our method can potentially be useful for such languages, due to its cost-effective nature. Next steps comprise the application of our methodology to a broader range of languages (including a variety of low resource languages) and adding datasets for more mental health tasks and from different social media platforms.
Limitations
Translation
Creating a multilingual dataset requires significant resources. In this work, we used a popular LLM like GPT3.5 to translate posts. Translating using only an LLM and not having an expert or native-language human resources introduces a small loss of information that in some cases affects the final results. Additionally, cultural differences are not “transferred" correctly when using machine translation.
Evaluation
Automatically assessing the performance of LLMs is inherently challenging. To evaluate performance, we search for the assigned label within the LLM’s output. If no label is detected, we default to assigning the minimum label—class 0 for the depression dataset and class 0 (not suicidal) for the suicide dataset.
Potential risks
The quality of publicly available datasets is crucial, especially in sensitive areas such as mental health care. The datasets we utilized, as well as the multilingual data we created for this study, must be handled with great caution. These datasets should serve as tools to support healthcare professionals and their research, rather than being used to diagnose patients directly.