### v2

# Building Multilingual Datasets for Predicting Mental Health Severity through LLMs: Prospects and Challenges

Abstract
Large Language Models (LLMs) are increasingly being integrated into various medical fields, including mental health support systems. However, there is a gap in research regarding the effectiveness of LLMs in non-English mental health support applications. To address this problem, we present a novel multilingual adaptation of widely-used mental health datasets, translated from English into six languages (e.g., Greek, Turkish, French, Portuguese, German, and Finnish). This dataset enables a comprehensive evaluation of LLM performance in detecting mental health conditions and assessing their severity across multiple languages. By experimenting with GPT and Llama, we observe considerable variability in performance across languages, despite being evaluated on the same translated dataset. This inconsistency underscores the complexities inherent in multilingual mental health support, where language-specific nuances and mental health data coverage can affect the accuracy of the models. Through comprehensive error analysis, we emphasize the risks of relying exclusively on LLMs in medical settings (e.g., their potential to contribute to misdiagnoses). Moreover, our proposed approach offers significant cost savings for multilingual tasks, presenting a major advantage for broad-scale implementation.
Building Multilingual Datasets for Predicting Mental Health Severity through LLMs: Prospects and Challenges
Konstantinos Skianis, A. Seza Doğruöz#, John Pavlopoulos Department of Computer Science and Engineering, University of Ioannina, Greece # LT3, IDLab, Universiteit Gent, Belgium Department of Informatics, Athens University of Economics and Business, Greece Archimedes/AthenaRC, Greece kskianis@cse.uoi.gr  as.dogruoz@ugent.be  annis@aueb.gr
1 Introduction
Warning: this paper contains text examples that are negative, depressive, or adverse.
Large language models (LLMs) (e.g., GPT-3, BERT, and their multilingual variants) are trained on vast amounts of text data and demonstrate an impressive ability to generate human-like text, perform complex language tasks Devlin et al. (2019); Brown et al. (2020). One of the most promising applications of LLMs is in the healthcare domain, where they have the potential to assist medical specialists with the detection, diagnosis, and treatment of health conditions of patients.
Mental health disorders (e.g., depression, anxiety, and post-traumatic stress disorder (PTSD)) are widespread across the globe, presenting major public health challenges World Health Organization (2021). Traditional diagnostic approaches typically involve (manual)-reported questionnaires, clinical interviews, and standardized assessments performed by trained professionals Kessler and Üstün (2004). However, these processes are also time-consuming and labor intensive, leading to delays in treatment and increased waiting times for the patients.
Therefore, leveraging large language models (LLMs) to detect mental health symptoms from textual data has presented a promising alternative. These models, for instance, have the ability to process vast amounts of textual content (e.g., social media posts, online forum discussions, personal narratives), and detect specific linguistic indicators related to mental health issues Guntuku et al. (2019); Chancellor et al. (2019). In this way, health professionals can be assisted with the early detection and intervention of mental health symptoms, reduce delays for patients whose symptoms might otherwise go unnoticed.
However, most studies about applying LLMs to mental health research are still in English and there is a need to explore this research area for other languages Raihan et al. (2024). Our study fills these gaps with the following research questions:
  • How accurately can LLMs predict the severity of mental health conditions in English-language through user-generated content?
  • When English-language mental health datasets are translated into other languages using LLMs, does the model maintain similar accuracy in predicting the severity of mental health conditions across these languages?
We answer these research questions with the following key contributions:
  • We create a novel multilingual dataset (covering English, Turkish, French, Portuguese, German, Greek, and Finnish) derived from user-generated social media content, which is automatically translated from English using an LLM. [^1] Code and some results are uploaded as supplementary material. The dataset will be made publicly available upon acceptance.
  • We perform an extensive evaluation of the effectiveness of an LLM in predicting the severity of mental health conditions across these different languages.
  • We provide an in-depth analysis of the model’s performance in each language, offering insights into how linguistic diversity affects the task and dataset-specific outcomes.
To the best of our knowledge, this is the first study to utilize LLMs for creating and evaluating multilingual adaptations of existing mental health NLP datasets, specifically designed to detect the severity of mental health conditions. Our study pushes forward the creation of inclusive and adaptable AI-powered tools and datasets for mental health diagnosis, emphasizing the critical need for approaches that account for cultural and linguistic diversity in mental health care, especially in non-English contexts with cost-efficient methods.
2 Related work
Early investigations into the use of NLP for mental health concentrated on identifying specific symptoms from textual data through machine learning methods. This often involved analyzing social media content, online health forum discussions, and other text-based sources to find indicators of mental health conditions like depression, anxiety and others De Choudhury and De (2014); Coppersmith et al. (2014); De Choudhury and De (2014); Coppersmith et al. (2015); Shing et al. (2018); Guntuku et al. (2019). Traditional approaches primarily relied on machine learning algorithms with hand-crafted features, which lacked the sophistication of current LLMs.
Multilingual Capabilities of LLMs
Although training and evaluation details are not always fully disclosed, LLMs are assumed to be capable of processing and generating text in multiple languages. For instance, the XLM-R model (trained on extensive multilingual data), has demonstrated strong cross-lingual performance. As noted by Conneau et al. (2020), XLM-R surpasses earlier models on various tasks, highlighting the benefits of leveraging large-scale multilingual data.
There are also examples of LLMs excelling in practical applications (e.g., translation and multilingual question answering). For instance, a model created by Xue et al. (2021), is pre-trained on a diverse set of languages and demonstrates strong capabilities in translating and answering questions in multiple languages. The model’s architecture and training methods allow it to generalize across languages effectively, providing robust performance even in languages with limited training data.
Figure 1: An illustration of our proposed methodology.
However, major obstacles still hinder the achievement of truly equitable performance across all languages Zhang et al. (2020). One of the primary challenges lies in low-resource languages, which suffer from a scarcity of digitally available text data. This often leads to under-performance by LLMs, producing less reliable outputs for such languages. Tackling this issue requires not only more inclusive and comprehensive data collection but also innovative approaches in transfer learning that can better leverage sparse resources Doğruöz and Sitaram (2022).
Moreover, language is intricately tied to cultural and societal frameworks, making it particularly difficult to capture the subtle, context-driven nuances that vary from one culture to another. This adds another layer of complexity to multilingual NLP tasks. Addressing these intricacies demands a twofold approach: developing more advanced, culturally-aware algorithms and incorporating datasets that are truly representative of linguistic and cultural diversity Doğruöz et al. (2023). Without these efforts, the full potential of multilingual LLMs will remain inaccessible to many, limiting their impact on global applications.
LLMs for Mental Health
LLMs excel at detecting mental health symptoms with greater precision by leveraging their advanced understanding of context and semantics. Notable examples include BioBERT Lee et al. (2020) and ClinicalBERT Huang et al. (2019), which have been pre-trained on biomedical literature and clinical notes. Additionally, models like MentalBERT Ji et al. (2022), DisorBERT Aragon et al. (2023), SuicidalBERT Anonymous , MentalLLM Xu et al. (2024) and MentalLlama Yang et al. (2023) are specifically trained on mental health-related social media data. Nevertheless, using LLMs for detecting mental health issues may introduce a number of advantages, as well as disadvantages De Choudhury et al. (2023). Additionally, exploiting social network posts represents a promising frontier in both artificial intelligence and mental health care Harrigian et al. (2020). LLMs (e.g., GPT-4 and BERT) can analyze vast amounts of textual data from social media platforms automatically to identify linguistic patterns associated with mental health conditions such as depression, anxiety, and suicidal ideation Ji et al. (2023).
Recently, a wide range of social network datasets focusing on mental health have become accessible Raihan et al. (2024). The authors compiled posts from platforms like Reddit and Twitter that pertain to conditions such as depression, PTSD, schizophrenia, and eating disorders. In addition, they fine-tuned several models using smaller, publicly available annotated mental health datasets to facilitate labeling for their newly introduced MentalHelp dataset. However, the dataset is limited to English-language posts, which constrains its application and hinders further research in multilingual contexts.
3 Proposed Methodology
We present a comprehensive methodology for leveraging LLMs and social media posts (in English), to translate them into multiple languages and predict mental health conditions accordingly.
First, we translate the social media posts across the proposed languages via an LLM. Then, these translations are fed into a prompt that asks the LLM to predict specific severity levels of mental health conditions. Finally, the predicted classes are compared with the ground truth labels and the evaluation metrics are computed.
Our goal is to propose a methodology that can be used to evaluate LLMs on multilingual detection of mental health conditions in languages where minimal relative linguistic resources are publicly available. An illustration of the proposed approach for evaluating LLMs for multilingual detection of mental health conditions is shown in Figure 1.
Dataset Reference Category #Classes #Instances Labels (#Support)
Dep-Sev Naseem et al. (2022) Depression 4 3553 Minimum (2587)
Mild (290)
Moderate (394)
Severe (282)
C-ssrs Gaur et al. (2019) Suicide 5 500 Supportive (108)
Indicator (99)
Ideation (171)
Behavior (77)
Attempt (45)
Table 1: Benchmark datasets used in our study.
4 Experiments
4.1 Datasets
In order to test the proposed methodology (above), we use the following datasets:
Dep-Severity
Introduced by Naseem et al. (2022), this dataset comprises posts from the social media platform Reddit, specifically addressing the varying levels of depression. These posts, all in English, are pre-labeled according to four severity levels: minimal, mild, moderate, and severe depression. Notably, the dataset is highly imbalanced, with the majority of posts classified under the minimal severity level.
C-SSRS
The C-SSRS (Columbia-Suicide Severity Rating Scale) dataset, created by Gaur et al. (2019), is a collection of structured data used to assess suicidal ideation and behavior. It is based on the Columbia-Suicide Severity Rating Scale, a widely used tool in clinical and research settings to evaluate the severity and frequency of suicidal thoughts, past attempts, and related risk factors. The dataset includes several variables (e.g., suicidal ideation intensity, past attempts, preparatory behaviors, and protective factors) allowing researchers and mental health professionals to identify high-risk individuals.
We deliberately chose multi-class datasets to increase the difficulty for the LLM, as binary classification tasks would have been simpler to solve. Table 1 provides a summary of key statistics for the two datasets utilized in our study.
Dataset Prediction prompt
Dep-Sev “Categorize the following text with 1 of the 4 depression severity levels (0: Minimum, 1: Mild, 2: Moderate, 3: Severe)”
C-ssrs “Categorize the content into the following classes based on the user’s mental state: 1. Supportive, 2. Indicator, 3. Ideation, 4. Behavior, 5. Attempt.”
Table 2: The used prediction prompts.
4.2 Settings
We use GPT3.5-turbo Brown et al. (2020), GPT-4o-mini via the OpenAI API and Llama3.1 locally to translate the posts and predict the labels. Recently introduced models, such as MentaLLaMA Yang et al. (2024), primarily focus on English, making them unsuitable for use in multilingual settings.The temperature parameter is set to 0, so the outcome is reproducible (regarding the translations and predictions). We experiment with no examples in the prompt (0-shot) and one example per class (1-shot). The task is solved as a classification problem, where we compare the predicted classes with the ground-truth ones, reporting Precision (Pr), Recall (Rec) and F1. We experiment with seven languages from major language families: Turkish (Altaic), French, Portuguese (Romance), English, German (Germanic), Finnish (Uralic), Greek (Hellenic). Table 2 shows the used prompts for predicting the severity levels.
English Turkish French Portuguese German Greek Finnish
Class Pr Rec F1 Pr Rec F1 Pr Rec F1 Pr Rec F1 Pr Rec F1 Pr Rec F1 Pr Rec F1
minimum 0.98 0.14 0.25 0.93 0.03 0.05 0.99 0.15 0.27 0.98 0.11 0.20 0.97 0.24 0.38 0.99 0.07 0.14 0.98 0.08 0.15
mild 0.04 0.15 0.07 0.02 0.01 0.01 0.05 0.17 0.08 0.04 0.07 0.05 0.08 0.32 0.12 0.04 0.17 0.06 0.05 0.19 0.08
moderate 0.13 0.22 0.17 0.10 0.62 0.17 0.15 0.39 0.21 0.12 0.28 0.17 0.16 0.42 0.23 0.14 0.55 0.23 0.12 0.31 0.17
severe 0.13 0.71 0.22 0.13 0.39 0.19 0.15 0.58 0.23 0.12 0.74 0.20 0.16 0.35 0.22 0.16 0.28 0.20 0.12 0.55 0.20
Macro avg 0.32 0.30 0.17 0.29 0.26 0.11 0.33 0.32 0.20 0.31 0.30 0.16 0.34 0.33 0.24 0.33 0.27 0.16 0.32 0.28 0.15
Table 3: GPT-3.5 with 0-shot learning on Dep-Severity, measuring Precision, Recall, and F1 per class per language. The last row shows the macro averages. The best F1 per class is shown in bold. The best macro-averaged F1 across languages is underlined.
English Turkish French Portuguese German Greek Finnish
Class Pr Rec F1 Pr Rec F1 Pr Rec F1 Pr Rec F1 Pr Rec F1 Pr Rec F1 Pr Rec F1
Supportive 0.27 0.13 0.18 0.40 0.02 0.04 0.30 0.19 0.24 0.36 0.04 0.07 0.38 0.27 0.31 0.39 0.35 0.37 0.29 0.09 0.14
Indicator 0.21 0.76 0.33 0.20 0.95 0.34 0.21 0.72 0.32 0.21 0.92 0.34 0.27 0.57 0.37 0.22 0.49 0.31 0.22 0.44 0.30
Ideation 0.45 0.26 0.33 0.25 0.05 0.08 0.40 0.21 0.28 0.30 0.09 0.14 0.40 0.49 0.44 0.39 0.42 0.40 0.36 0.56 0.44
Behavior 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.50 0.03 0.05 0.00 0.00 0.00 0.00 0.00 0.00
Attempt 0.00 0.00 0.00 0.25 0.02 0.04 0.00 0.00 0.00 1.00 0.02 0.04 1.00 0.04 0.09 0.00 0.00 0.00 0.00 0.00 0.00
Macro avg 0.19 0.23 0.17 0.22 0.21 0.10 0.18 0.22 0.17 0.37 0.21 0.12 0.51 0.28 0.25 0.20 0.25 0.22 0.17 0.22 0.18
Table 4: GPT-3.5-turbo with 0-shot learning on C-ssrs, measuring Precision, Recall, and F1 per class per language. The last row shows the macro averages. The best F1 per class is shown in bold. The best average F1 across languages is underlined.
English Turkish French Portuguese German Greek Finnish
Class Pr Rec F1 Pr Rec F1 Pr Rec F1 Pr Rec F1 Pr Rec F1 Pr Rec F1 Pr Rec F1
minimum 0.97 0.32 0.48 0.97 0.18 0.30 0.97 0.30 0.46 0.97 0.31 0.47 0.96 0.36 0.52 0.98 0.26 0.41 0.97 0.26 0.40
mild 0.07 0.26 0.11 0.04 0.18 0.07 0.05 0.14 0.07 0.07 0.22 0.11 0.06 0.14 0.08 0.05 0.13 0.07 0.05 0.17 0.07
moderate 0.18 0.41 0.25 0.15 0.42 0.23 0.16 0.50 0.24 0.16 0.41 0.23 0.15 0.42 0.22 0.15 0.52 0.23 0.15 0.48 0.23
severe 0.19 0.43 0.26 0.17 0.48 0.25 0.18 0.44 0.26 0.17 0.49 0.25 0.18 0.48 0.26 0.17 0.40 0.24 0.18 0.40 0.25
Macro avg 0.35 0.35 0.27 0.33 0.31 0.21 0.34 0.34 0.26 0.34 0.36 0.27 0.34 0.35 0.27 0.34 0.33 0.24 0.34 0.33 0.24
Table 5: GPT-4o-mini with 0-shot learning on Dep-Severity, measuring Precision, Recall, and F1 per class per language. The last row shows the macro averages. The best F1 per class is shown in bold. The best macro-averaged F1 across languages is underlined.
English Turkish French Portuguese German Greek Finnish
Class Pr Rec F1 Pr Rec F1 Pr Rec F1 Pr Rec F1 Pr Rec F1 Pr Rec F1 Pr Rec F1
Supportive 0.40 0.65 0.50 0.40 0.69 0.51 0.43 0.60 0.50 0.49 0.57 0.53 0.44 0.56 0.50 0.45 0.54 0.49 0.39 0.70 0.50
Indicator 0.26 0.52 0.34 0.27 0.36 0.31 0.27 0.44 0.33 0.29 0.55 0.38 0.24 0.46 0.32 0.27 0.53 0.36 0.21 0.29 0.25
Ideation 0.48 0.20 0.29 0.46 0.35 0.40 0.46 0.32 0.37 0.43 0.26 0.33 0.42 0.36 0.39 0.46 0.46 0.46 0.44 0.37 0.40
Behavior 1.00 0.01 0.03 0.31 0.06 0.11 0.67 0.03 0.05 0.50 0.01 0.03 0.00 0.00 0.00 0.00 0.00 0.00 0.50 0.01 0.03
Attempt 0.39 0.47 0.42 0.30 0.22 0.26 0.31 0.44 0.37 0.34 0.62 0.44 0.42 0.22 0.29 0.57 0.09 0.15 0.35 0.13 0.19
Macro avg 0.51 0.37 0.32 0.35 0.34 0.32 0.43 0.37 0.33 0.41 0.40 0.34 0.30 0.32 0.30 0.35 0.32 0.29 0.38 0.30 0.27
Table 6: GPT-4o-mini with 0-shot learning on C-ssrs, measuring Precision, Recall, and F1 per class per language. The last row shows the macro averages. The best F1 per class is shown in bold. The best average F1 across languages is underlined.
English French Portuguese German
Class Pr Rec F1 Pr Rec F1 Pr Rec F1 Pr Rec F1
minimum 0.88 0.12 0.22 0.91 0.23 0.37 0.95 0.18 0.30 0.95 0.25 0.40
mild 0.04 0.20 0.07 0.04 0.13 0.07 0.05 0.14 0.08 0.05 0.15 0.08
moderate 0.15 0.60 0.25 0.16 0.58 0.24 0.14 0.76 0.24 0.14 0.41 0.21
severe 0.18 0.22 0.20 0.12 0.26 0.17 0.20 0.11 0.14 0.13 0.41 0.20
Macro avg 0.32 0.29 0.19 0.31 0.30 0.21 0.34 0.30 0.19 0.32 0.31 0.22
Table 7: Llama3.1 with 0-shot learning on Dep-Severity, measuring Precision, Recall, and F1 per class per language. The last row shows the macro averages. The best F1 per class is shown in bold. The best average F1 across languages is underlined.
English Turkish French Portuguese German Greek Finnish
Class Pr Rec F1 Pr Rec F1 Pr Rec F1 Pr Rec F1 Pr Rec F1 Pr Rec F1 Pr Rec F1
minimum 0.90 0.14 0.25 0.76 0.23 0.36 0.91 0.11 0.19 0.90 0.11 0.20 0.76 0.75 0.76 0.79 0.02 0.05 0.68 0.16 0.25
mild 0.03 0.06 0.04 0.07 0.10 0.08 0.04 0.04 0.04 0.05 0.14 0.07 0.07 0.08 0.07 0.05 0.10 0.06 0.03 0.08 0.05
moderate 0.14 0.65 0.23 0.11 0.29 0.16 0.13 0.44 0.20 0.15 0.51 0.23 0.16 0.23 0.19 0.14 0.26 0.18 0.13 0.39 0.19
severe 0.15 0.35 0.21 0.09 0.40 0.14 0.11 0.61 0.19 0.13 0.48 0.20 0.16 0.07 0.09 0.10 0.73 0.17 0.10 0.37 0.16
Macro avg 0.30 0.30 0.18 0.26 0.26 0.18 0.30 0.30 0.16 0.31 0.31 0.18 0.29 0.28 0.28 0.27 0.28 0.12 0.23 0.25 0.16
Table 8: GPT-3.5 with 1-shot learning on Dep-Severity by measuring Precision, Recall, and F1 per class per language. The last row shows macro averages for precision, recall, and F1. Best F1 per class is shown in bold. Best average F1 across languages is underlined.
4.3 Preliminary Prompting
Before presenting any posts to our LLM, whether for translation or classification, we first inquired how it would approach categorizing them into varying levels of depression severity. The LLM explained that it would begin by identifying specific language cues linked to depressive states, such as:
  • Recurrent negative emotions like sadness or a sense of hopelessness.
  • Self-critical thoughts or feelings of worthlessness.
  • Expressions reflecting isolation or withdrawal from social interactions.
  • Observable shifts in behavior or routines (e.g., disrupted sleep or changes in appetite).
  • Mentions of emotional suffering or psychological distress.
Furthermore, the LLM would attempt to adapt these signals to the four depression severity levels in the context of social media content. For instance:
  • Level 1 (Minimal): Posts that show infrequent or minor expressions of sadness or frustration.
  • Level 2 (Mild): Posts where negative emotions or notable behavioral shifts are more consistent.
  • Level 3 (Moderate): Posts indicating substantial interference with daily life or clear markers of emotional distress.
  • Level 4 (Severe): Posts showing extreme emotional distress, risk of self-harm, or complete social disengagement.
From this, we can infer that the LLM anticipates posts containing broad indicators of depressive symptoms.
Similarly, for the second dataset, we expect the LLM to detect various linguistic and contextual features in the social media posts to classify them. Some key aspects that it considers per class are:
  1. 1.
    Supportive (Encouragement and Positive Reinforcement). Some key linguistic features that may appear include expressions of empathy, encouragement, positivity and use of supportive phrases (e.g., ”You are not alone”, ”Stay strong”, ”I’m here for you”, ”You matter”).
  2. 2.
    Indication (Signs of Distress or Risk). Linguistic features may include the use of worrying expressions (e.g., ”I feel lost”, ”I don’t know what to do”, ”Nothing makes sense anymore”, ”I’m so tired”; indirect mentions of mental health struggles: ”I’m always exhausted”, ”Everything is meaningless”, ”I just want it to stop”).
  3. 3.
    Ideation (Expressions of Suicidal Thoughts). Key linguistic features may include direct expressions of suicidal thoughts (e.g., ”I wish I didn’t exist”, ”I want to end it all”, ”I don’t see the point of living”; questions about death and suicide: ”What happens after we die?”, ”Is there peace after death?”; use of hopelessness in an extreme form: ”Nothing will ever change”, ”I just want to disappear”).
  4. 4.
    Behavior (Actions Related to Suicide). Key linguistic features include descriptions of planning or preparation (e.g., ”I’ve been researching painless ways to go”, ”I wrote my letter today”, ”This might be my last post”, ”I have the pills ready”; mentions of giving away possessions or saying goodbye: ”I just want to thank everyone for being in my life”).
  5. 5.
    Attempt (Reports of Actual Suicide Attempts). Key linguistic features include explicit mentions of self-harm or suicide attempts (e.g., ”I tried last night but failed”, ”I was in the hospital after overdosing”, ”I attempted but someone stopped me”, ”I survived but I don’t know if I should be happy about it”) and descriptions of medical intervention (e.g., ”They pumped my stomach”, ”They found me just in time”, ”I’m in the ER right now”).
4.4 Experimental results
The results are organized and presented in the following subsections: classification in the source language, classification in the target language, classification across languages, and further analysis of LLMs performance with multilingual examples.
Source language
Initially, we experimented with the data in their source language (English), to set the baseline performance. In other words, no translation step has been performed at this stage. When using GPT-3.5, the first column in Tables 3 and 4 presents the results on depression and suicide detection tasks respectively. The best F1 in both tasks is achieved for the lowest severity/indication (minimum depression and supportive for suicide). In Dep-Severity, the best performance is achieved for the two edges of lowest (F1=0.48) and highest severity (F1=0.26). In C-ssrs, on the other hand, F1 drops from 0.18 to 0.00. When using the superior GPT-4o-mini (Tables 5 and 6 ), F1 increases to 0.27 and to 0.32, respectively. Detecting specific levels of suicidal ideation is harder than in the case of depression, as they may be considered less distinct. Therefore, it could be probably more challenging for the LLM to distinguish between them, also shown by the results for some classes.
Target languages
GPT-3.5 has a varying performance when operating on content translated from English to target languages. In Dep-Severity (Table 3), performance drops in Turkish (-6), Portuguese (-1), Greek (-1), and Finish (-2). On the other hand, the performance increases in French (+3) and German (+7). In C-ssrs (Table 4), performance drops in all target languages, with the drop for Turkish (-7) being the most considerable one. This can be explained by the limited number of Turkish resources in health, especially in the domain of mental health Çöltekin et al. (2023).
GPT-4o-mini in Dep-Severity (Table 5) performs equally well in Portuguese and German (as in English). However, this is not true for the rest of the target languages. In C-ssrs, performance remains similar for most target languages. However, we observe again the phenomenon of better performance for a target language, namely for French (+1) and Portuguese (+2), as was the case for GPT-3.5 in German. In general, GPT-4o-mini is much more stable and consistent for correct prediction.
Classification across languages
Figures 2 and 3 present F1 in boxplots, summarising the performance of GPT-3.5 across languages per task. In Dep-Severity, the lowest (minimum) severity class is characterized by a wide range, from 0.15 to 0.25. The whiskers are also wide, from 0.05 (Turkish) to 0.38 (German). Mild severity is mishandled across languages while for the classes of moderate and severe, the model achieves a relatively high score across languages.
Figure 2: F1 of GPT-3.5 across languages (vertically) per class (box) for Dep-Severity.
Figure 3: F1 of GPT-4o-mini across languages (vertically) per class (box) for C-ssrs.
In C-ssrs, the best performance is achieved for the “Supportive” category. This is also the second major class, while slightly worse performance is achieved for the other classes. For all classes, apart from “Attempt” the performance is similarly distributed across languages.
From zero to single shot
Table 8 presents the results when 1-shot learning is employed. In English, performance remains similarly low in Dep-Severity. Again, performance is better for a target language (German), compared to the performance for the source language. In general, we observe an increase going from 0-shot to 1-shot. Better performance could be observed if more examples were provided. However, this would also increase the cost of the experiments by approx. doubling the size of the prompt.
Experimenting with Llama-3.1
Table 7 shows the performance of Llama-3.1 in Dep-Severity using 0-shot, for the languages it was available for. Although the model performs better in English compared to GPT-3.5 (both 0- and 1-shot), it is slightly worse in target languages. Notably, the best performance is not achieved in English.
Back-translation
To assess the effect of translation, we translated the already translated texts back to English and repeated the experiments for GPT-3.5-turbo with 0-shot on C-ssrs. Table 9 shows that the prediction increased when translating back to English. The highest gain (from 0.17; see Table 4) is observed when translating to Portuguese or Finnish and back (+0.05) while +0.03 is reported for Turkish and +0.02 for the rest. Hence, we observe that translating before predicting assists the prediction performance of the LLM. We present some representative examples in the Appendix.
Turkish French Portuguese German Greek Finnish
Class Pr Rec F1 Pr Rec F1 Pr Rec F1 Pr Rec F1 Pr Rec F1 Pr Rec F1
Supportive 0.43 0.33 0.38 0.37 0.26 0.30 0.44 0.34 0.39 0.37 0.31 0.34 0.36 0.19 0.24 0.40 0.34 0.37
Indicator 0.22 0.77 0.35 0.22 0.75 0.34 0.23 0.76 0.36 0.20 0.64 0.31 0.22 0.79 0.34 0.23 0.74 0.35
Ideation 0.43 0.19 0.27 0.42 0.20 0.27 0.40 0.22 0.28 0.41 0.22 0.29 0.45 0.22 0.29 0.41 0.21 0.28
Behavior 1.00 0.01 0.03 0.67 0.03 0.05 0.00 0.00 0.00 1.00 0.01 0.03 0.50 0.01 0.03 0.00 0.00 0.00
Attempt 0.00 0.00 0.00 0.00 0.00 0.00 1.00 0.04 0.09 0.00 0.00 0.00 0.50 0.02 0.04 1.00 0.04 0.09
Macro avg 0.42 0.26 0.20 0.33 0.25 0.19 0.42 0.27 0.22 0.40 0.24 0.19 0.40 0.24 0.19 0.41 0.27 0.22
Table 9: Back-translation and prediction with GPT-3.5-turbo with 0-shot learning on C-ssrs, measuring Precision, Recall, and F1 per class per language. The last row shows the macro averages. The best F1 per class is shown in bold. The best average F1 across languages is underlined.
Further analysis with examples
Figure 4: An example of translation (English to Greek) from the C-ssrs dataset with GPT-4o-mini.
Lastly, Figure 4 presents two instances of the suicide dataset. Regarding the first example, while the translation from English to Greek appears to be correct in terms of conveying the same message, the English text is classified as ”Supportive”, while the Greek text is classified incorrectly as ”Indicator”. In this case, GPT could interpret the context differently due to the nuances in the languages. In Greek, the model might be weighing other parts of the text more heavily, leading to a classification of ”Indicator”. This could be due to differences in the training data, the intricacies of the language model, or contextual emphasis.
For the second example, we observe that the translation is accurate and conveys the same message as the original English text. Nevertheless, the English text was incorrectly classified as level 2 (indicator). However, the Greek translation was correctly classified as level 3 (ideation). Here, we see that detecting different levels is inherently difficult. The original English post could be mispredicted as it is quite small and generic.
4.5 Resources and cost of experiments
The total cost for running the GPT3.5 experiments via its API was less than \$30, showing the minimal resources required to conduct them, without the need for expensive GPU infrastructure or fine-tuning. For example, we experimented with MentaLLaMA-chat-7B Yang et al. (2024), which is a 27GB model. Its size makes it prohibitive when limited resources are available. Eventually, we ended up using the quantized version, which did not output meaningful results. Our methodology for utilizing resources is potentially promising for extending the medical data sets (mostly in English) into other languages as well.
5 Discussion
Despite the hype of using LLMs in mental healthcare for commercial applications, there is an increasing need for more research to understand the pitfalls and dangers of using LLMs across languages, diseases, and patient populations for mental health care. Considering the variation in the performance of the LLMs across languages, there is a need for utmost precautions not to rely on them solely in any healthcare setting. As stated by Stade et al. (2024), mental health diagnosis should never be left to automatic systems; human professionals are critical due to the possibility of errors and misdiagnoses, as illustrated above.
LLMs exhibit considerable variability in performance across languages, despite being evaluated on the same translated dataset. This inconsistency underscores the complexities inherent in multilingual mental health support, where language-specific nuances and mental health data coverage can effect the accuracy of the models. We observe performance gains in both tasks with stronger models and by translating to a specific target language (e.g., in German, for depression severity) before detection. On the other hand, translating to Turkish yields the worst performance in both tasks, which could be explained by the limited number of Turkish resources in health, especially in the domain of mental health Çöltekin et al. (2023).
6 Conclusion
We presented a novel multilingual dataset designed to evaluate the ability of LLMs to predict the severity of mental health conditions in user-generated posts. Using English as the source language, we translated the posts to French, German, Turkish, Greek, Finnish, and Portuguese, using LLM-based machine translation. Then, we prompted the LLM to predict the severity of two mental health conditions across all languages. Our results indicate that the LLM performance varies depending on the language. Also, the best performance is not necessarily achieved in the source (English) language. Tested on low-resource languages (i.e., Greek and Turkish on medical domain), our method is useful for other languages due to its cost-effective nature. Next steps comprise the application of our methodology to a broader range of languages (including a variety of low-resource languages) and adding datasets for more mental health tasks.
Limitations
Translation
Creating a multilingual dataset requires significant resources. In this work, we used a popular LLM like GPT3.5 to translate posts. Translating using only an LLM and not having an expert or native-language human resources introduces a small loss of information that in some cases affects the final results. Additionally, cultural differences are not “transferred” correctly when using machine translation.
Evaluation
Automatically assessing the performance of LLMs is inherently challenging. To evaluate performance, we search for the assigned label within the LLM’s output. If no label is detected, we assign the minimum label—class 0 for the depression dataset and class 2 for the suicide dataset (Indicator).
Potential risks
The quality of publicly available datasets is crucial, especially in sensitive areas such as mental health care. The datasets we utilized, as well as the multilingual data we created for this study, must be handled with great caution. These datasets should serve as tools to support healthcare professionals and their research, rather than being used to diagnose patients directly.
Appendix A Appendix
Next, we present some examples of misprediction when using the original post, translated and back-translated text. All examples come from the C-ssrs dataset.
A.1 Translation mispredicted, back-translation ok
Original post: “Im sure that no ones lying to you. A man I know just succeeded in committing Suicide and I can assure you that he was a ghost of himself, consumed by major Mental Depression for months. Everyone cares, and everyone wishes he hadnt done it. I hope you can find a way to carry on, and I hope that if you give it some time you can find healing and happiness in your life.”
Ground-truth: 2. Indicator
Prediction: 2. Indicator
Here the original post was correctly classified as “2. Indicator”.
Finnish translation: “Olen varma, ettei kukaan valehtele sinulle. Mies, jonka tunnen, onnistui juuri tekemään itsemurhan, ja voin vakuuttaa, että hän oli varjo entisestään, syötyään pahoista mielenterveysongelmista kuukausia. Kaikki välittävät, ja kaikki toivovat, ettei hän olisi tehnyt sitä. Toivon, että löydät tavan jatkaa eteenpäin, ja toivon, että jos annat sille aikaa, voit löytää parantumista ja onnea elämääsi.”
Prediction: 3. Ideation
Back-translation to English: “I am sure that no one is lying to you. The man I know just managed to commit suicide, and I can assure you that he was a shadow of his former self, having suffered from severe mental health issues for months. Everyone cares, and everyone wishes he hadn’t done it. I hope you find a way to move forward, and I hope that if you give it time, you can find healing and happiness in your life.”
Prediction with back-translation: 2. Indicator
In this example we observe that the back-translation helped to correct the prediction.
A.2 Translation ok, back-translation misprediction
Original post: “Dont feel guilty. Better than doing nothing, get someone(his family or a friend) to show him that he is cared for. If you get a chance you could go personally and help him. For now keep him encouraged. Try to cheer him up and distract him from these thoughts. I understand how you feel. I also get such thoughts. If you need a friend, Ill be there for you(you can PM me). Dont waste yourself. You are a blessing to this world.Your family will feel Tired sad. BTW, its better you try taking that anti-depressing medication. Please throw away the gun. True. Everyone has a guardian angel.”
Ground-truth: 1. Supportive
Prediction: 2. Indicator
Here the original post was misclassified as “2. Indicator” while it belongs to “1. Supportive”.
[…]
Πρεδιςτιον ωιτη Γρεεϰ τρανςλατιον: 1. Συππορτιε
Βαςϰ-τρανςλατιον το Ενγλιςη: ``Δο νοτ φεελ γυιλτψ. Ινςτεαδ οφ δοινγ νοτηινγ, φινδ ςομεονε (τηειρ φαμιλψ ορ α φριενδ) το ςηοω τηεμ τηατ ψου ςαρε. Ιφ ψου ηαε τηε οππορτυνιτψ, ψου ςαν γο ιν περςον ανδ ηελπ τηεμ. Φορ νοω, ϰεεπ τηεμ ενςουραγεδ. Τρψ το λιφτ τηειρ ςπιριτς ανδ διςτραςτ τηεμ φρομ τηοςε τηουγητς. Ι υνδερςτανδ ηοω ψου φεελ. Ι αλςο ηαε ςυςη τηουγητς. Ιφ ψου νεεδ α φριενδ, Ι ωιλλ βε ηερε φορ ψου (ψου ςαν μεςςαγε με). Δο νοτ ωαςτε ψουρςελφ. Ψου αρε α βλεςςινγ το τηις ωορλδ. Ψουρ φαμιλψ ωιλλ φεελ τιρεδ ανδ ςαδ. Αλςο, ιτ ις βεττερ το τρψ ταϰινγ τηις αντιδεπρεςςαντ μεδιςατιον. Πλεαςε τηροω τηε γυν αωαψ. Τρυε. Εερψ περςον ηας α γυαρδιαν ανγελ.᾽
Πρεδιςτιον ωιτη βαςϰ-τρανςλατεδ: 2. Ινδιςατορ.
Ωε οβςερε τηατ τηε τεξτ ωας ινςορρεςτλψ πρεδιςτεδ, ωηιλε τηε Γρεεϰ τρανςλατιον ωας πρεδιςτεδ ςορρεςτλψ.